OpenStudy (anonymous):

Write a quadratic equation in standard form if the zeros are ±7i.

OpenStudy (anonymous):

(x+7i)(x-7i)

OpenStudy (anonymous):

Zero means x=that. x=7i x=-7i To make these into factors, we get them equal to 0. x+7i=0 x-7i=0 We can put these into an equation now. (x+7i)(x-7i)=0 Expansion of complex conjugates is easy: (a+bi)(a-bi)=a^2+b^2 x^2+49=0

OpenStudy (anonymous):

If you want the work for expansion: (x+7i)(x-7i) x(x-7i) +7i(x-7i) x^2-7xi +7xi -7i*7i x^2+0xi-49i^2 i^2=sqrt(-1)^2=-1, so x^2-49*-1 x^2+49

OpenStudy (anonymous):

Finally get this now, thanks so much! This was a big help! =)